18.57 - Applications of Electrostatics Puntos:1 Calculate the angular velocity ω of an electron orbiting a proton in the hydrogen atom, given the radius of the orbit is 5.300 x 10 11 m. You may assume that the proton is stationary and the centripetal force is supplied by Coulomb attraction. 1.30 10418rad/s Enviar Respuesta Incorrecto. Tries 3/5 Intentos Anteriores
Answer #1

Solution)

We know, F= ma = Kq1q2/r^2

So a = Kq1q2/(mr^2)

a = (9e^9 * 1.6 e^-19 *1.6 e -19)/(9.11 e-31 * 0.53 e -10* 0.53 e-10)

So, a = 9*10^ 22 m/s^2

Now, we know, a = r W^2

so W^2 = a/r = 9.e 22/(0.53 *10^ -10)

W^2 = 1.698*10^33

W = 4.12 *10^16 rad/s (Ans)

========

Good luck!:)